OpenStudy (darkbluechocobo):

After how many seconds would Black Dragon and Wild Thing have the same amount of fuel left in their tanks? Round your answer to the nearest hundredth of a second. (Enter only the number in the form 0.123) Hint: Select the two relevant equations from the last problem and solve them for the variable s.

OpenStudy (anonymous):

How much fuel does Black Dragon and Wild Thing have at the beginning?

OpenStudy (darkbluechocobo):

@ranga Hey was just wondering if you can help me

OpenStudy (darkbluechocobo):

well in the past equations this is what it looked like I did them

OpenStudy (darkbluechocobo):

Does anyone know :/ I really need help

OpenStudy (anonymous):

Black Dragon Equation: f(t)=18-2.33t Wild Thing Equation: f(t)=20-3.06t So the equation you would use to get the answer is y=20-3.06t y=-18+2.33t If we set the two functions equal to each other... 20 - 3.06t = 18 - 2.33t 2 = 3.06t - 2.33t 2 = 0.73t 2/0.73 = t t = 2.74 sec

OpenStudy (darkbluechocobo):

Thank you :3

OpenStudy (anonymous):

Np